OpenStudy (amber15):

add equation (5x^8 -5z^4+5z^2+3)+(3z^6+4 z^4-9z) simplify answer to sum

OpenStudy (bakonloverk):

Do you know how to add polynomials?

OpenStudy (anonymous):

Let's simplify step-by-step. 5x8−5z4+5z2+3+3z6+4z4−9z =5x8+−5z4+5z2+3+3z6+4z4+−9z Combine Like Terms: =5x8+−5z4+5z2+3+3z6+4z4+−9z =(5x8)+(3z6)+(−5z4+4z4)+(5z2)+(−9z)+(3) =5x8+3z6+−z4+5z2+−9z+3 Answer: =5x8+3z6−z4+5z2−9z+3